# The 2010 Census Confidentiality Protections Failed, Here’s How and Why

### v1

@@ 4. Reconstruction Methodology
We define database reconstruction as any attempt to re-create the record-level image of the database from which a set of published query results or tabulations were originally calculated; in this case, that is the confidential HDF. [^14] See [garfinkel2018a, ] for a longer discussion of how to understand database reconstruction. Database reconstruction attempts to reverse-engineer the confidential HDF records that were the input data used in a tabulation system with the goal of making the reconstruction as close as possible to these confidential data. We note that the reconstruction described here is not the most powerful feasible reconstruction. We used only a subset of the SF1 tables, we made no attempt to reconstruct households or characteristics of the householder (adult respondent providing the census data on CUF), and we did not use statistical modeling to improve the reconstruction (e.g., in blocks where multiple different reconstructions were consistent with the published tables, we did not use statistical methods to identify the most likely reconstruction). Even without these enhancements, we show that the disclosure avoidance methods of the 2010 Census are highly susceptible to attack.
@@ 4.2. A Simplified Algebraic Representation of the Reconstruction Problem
where $\mathbf{c}$ is $K\times 1$ column vector of the $K$ statistics extracted from SF1 for a given reconstruction, $\mathbf{A}$ is the matrix for computing those SF1 tabulations from a contingency table vector $\mathbf{x}$, and $\mathbb{Z}^{+}$ is the set of non-negative integers. Each row of $\mathbf{A}$ and corresponding component of $\mathbf{c}$, therefore, represent the formula for a single statistic and its realized value in SF1, respectively.
There are fewer statistics published per block than points in the sample space ($|\chi|>K$), meaning that a unique solution to equation (1) is not always guaranteed. However, many low-population blocks have published tables that contain a large number of entries whose values are zero. This eliminates many candidate solutions, and often there is only one unique non-negative, integer-valued solution to equation (1). There is always at least one solution because SF1 was tabulated from a single real database with the schema encoded in $\chi$ (i.e., the HDF). The commercial software GurobiTM has a robust toolkit for solving IPs and related mixed integer linear programs (MILP) to produce solutions to problems such as these. See Section 4.3 for details. When multiple solutions exist, GurobiTM will pick one. In principle, statistical modeling could be used to select among candidate solutions to improve reconstruction quality, but we did not do this.
Given a solution $\hat{x}$ to equation (1), another question of interest is how different $\hat{x}$ could be from other possible solutions. If $\hat{x}$ is the only solution to equation (1), then it represents an exact reconstruction of the microdata used to tabulate SF1, namely HDF, with certainty. Moreover, if the uniqueness of the solution can be determined from the published inputs to equation (1), then an attacker knows that those reconstructed records were present in HDF with certainty. To examine how often the solution $\hat{x}$ was strongly constrained, we designed an algorithm for measuring solution variability by building a second IP model. Its goal is to find a second solution, $\tilde{x}$, that maximizes the $\text{L}_{1}$ distance to $\hat{x}$. This allows an attacker, using only public information, to determine the maximum number of records in the reconstruction that could be incorrect. This problem is described in Section 4.5.
@@ 4.3. The Integer Programming Version of the Reconstruction Model
AIAN =$=\{\text{American Indian or Alaskan Native}=1,\text{Not American Indian or Alaskan Native}=0\}$
@@
SEX =$=\{\text{Male}=0,\text{Female}=1\}$
AGE =$=\{0,1,2,\dots,110\}.$
A potential record corresponds to a setting of the block $b$ along with settings for the demographic attributes (w, bl, aian, asian, nhopi, sor, hisp, a, s). Here the lowercase italic letters index the permitted values of the uppercase feature set of the same name (e.g. $aian$ represents a value of the feature AIAN). We create one variable for the IP for each potential record. Since we do not know, a priori, how many records exist with the same demographic type and block, we must create multiple variables $\mathcal{B}_{(i,b,w,bl,aian,asian,nhopi,sor,hisp,a,s)}$ for $i=0,1,2,\dots$. We use Table P12 from SF1 (age group by sex for each census block) to obtain an upper bound on the index $i$ for each block and combination of demographics. For example, for the demographic type of a 22-year-old Asian Hispanic female in a block $b$, suppose Table P12 for that block indicates that there are 50 females in the age group 22-24 (See Table 4 for the age groupings in P12). This is an upper bound on the number of potential records for 22-year-old Asian Hispanic females in the block. Thus, we create the 50 variables $\mathcal{B}_{(i,b,w=0,bl=0,aian=0,asian=1,nhopi=0,sor=0,hisp=1,a=22,s=1)}$ for $i=0,\dots,49$. These variables are binary, with a value of 1 indicating the presence of the potential record in a candidate microdata reconstruction. Once the binary variables are assigned values, the summation
@@
  • For every block $b$, age $a$, sex $s$, and value of the race/ethnicity variables w, bl, aian, asian, nhopi, sor, hisp, we define the binary variables:
@@
where the summation over $i$ uses the upper bound on the number of $\mathcal{B}_{(\cdot)}$ variables in equation (3) (i.e., $i=0,\dots,c^{\text{P12}}_{b,\text{AGEBIN$\_$P12}(a),s}-1$).
@@
We use the notational shorthand $\mathcal{T}_{t}[tabname,\bigcup_{b^{\prime}\in\text{B}_{t}}\text{rHDF}_{b,t}(b^{\prime})]=c_{t}^{tabname}$ to represent all such constraints created by SF1 tract-level table $tabname$ (e.g., PCT12I) for tract $t$. Specifically, the notation $\text{rHDF}_{b,t}(b^{\prime})$ refers to the optimization variables for records in a block $b^{\prime}$ when performing the $\text{rHDF}_{b,t}$ reconstruction, the notation $\bigcup_{b^{\prime}\in\text{B}_{t}}\text{rHDF}_{b,t}(b^{\prime})$ indicates that we form the constraints using the optimization variables for all blocks $b^{\prime}$ that belong inside tract $t$ (i.e., $\forall b^{\prime}\in\text{B}_{t}$). For example, in the case of SF1 Table PCT12I, this means an application of equation (4.3) for each age bin $z$ and sex $s$ for the record variables in the tract. The tract-level tables used are listed in Panel B of Table 2. Similarly, we use the shorthand $\mathcal{T}_{b}[tabname,\text{rHDF}_{b,t}(b^{\prime})]=c_{b^{\prime}}^{tabname}$ for the block-level constraints in block $b^{\prime}$ created by SF1 block-level table $tabname$ (Panel A of Table 2).
@@
The objective function “max 0” indicates that any feasible solution that satisfies the constraints can be returned (i.e., if there are multiple candidate solutions, no statistical modeling or maximum entropy is used, and GurobiTM picks a solution arbitrarily). The optimization variables are the $\mathcal{B}_{(\cdot)}$ defined in equation (3). Once we have a feasible solution, for every $\mathcal{B}_{(\cdot)}$ that is set to 1, we add a corresponding record for that block and demographic type to the reconstructed dataset. For $\text{rHDF}_{b,t}$, we also recode the feature age into agebin because our matching algorithms require access to both variables. See Section 5. Thus, upon completion, the IP from equation (4) yields a reconstructed version of the HDF, called $\text{rHDF}_{b,t}$, that contains one record for each person in the 2010 Census with the features indicated in row $\text{rHDF}_{b,t}$ of Table 1.161616 Since SF1 always groups all persons 100 years of age or older into the bins: “100-104 years”, “105-109 years”, and “110 years and over” or uses even coarser age groups, the IP can never resolve the feature age more precisely than these bins. The IP still has variables for individual ages 100, …, 110 and so Gurobi will arbitrarily choose a specific age within those age bins. This means that there is inherent solution variability in the ages of the oldest sub-populations.
Next, we present the IP for the reconstruction $\text{rHDF}_{b}$, which uses only the SF1 block-level tables shown in Panel A of Table 2. Block-level tables use several age binning schemes, but their intersection is not exact age. Instead, their intersection is the age grouping shown in Table 5, which has 38 age bins. Since this is the most fine-grained age resolution that can possibly be obtained from block-level tables, $\text{rHDF}_{b}$ contains age reconstructed up to this 38-age binning (the feature that represents this age grouping is called agebin). To simplify the implementation, the IP uses the $\mathcal{B}_{(\cdot)}$ optimization variables from equation (3). This means that the solution provides single-year ages, but after the reconstructed microdata are created, we recode age into agebin as defined in Table 5. The IP for the block-level reconstruction is similar to the tract-level reconstruction. For each block $b^{\prime}$, solve
@@
4.4. Implementation Notes
The IP formulation of the reconstruction problem is mathematically equivalent to the simplified model in Section 4.1. The only difference is that the optimization variables $\mathcal{B}_{(\cdot)}$ in the IP implemented in our code correspond to potential records, while the vector $\mathbf{x}$ in equation (1) contains counts from the fully saturated contingency table defined on the sample space $\chi$. The code in the replication archive implements the IP in equations (4) and (5), and these equations are useful for reading the software implementation in the replication archive, whereas the histogram representation in equation (1) is useful for understanding the high-level structure of the problem. For reconstruction outcomes, only run-time, not the space of feasible solutions, is affected by the choice of which representation to implement. It is not obvious a priori which representation should solve more quickly. We have direct experience solving the IP representation, and we found it to consistently solve very quickly at default GurobiTM settings for the set of tables we used in the reconstruction. Equation (1) is succinct, easy to represent in any matrix programming language that implements sparse matrix storage and MILP solvers, and yields a model with considerably fewer variables. On the other hand, the IP representation, while it produces less succinct models, uses only binary variables in the solution set, rather than non-negative integers, and binary variables are usually processed more efficiently in modern MILP solvers. Our discussion switches between the two representations to permit clarity of expression (equation (1)) versus fidelity to the details of our implemented reconstruction code (equations (4) and (5)).
@@ 4.5. Solution Variability
where feasible solutions use equation (4) when $g$ is a tract or equation (5) when $g$ is a block. Note that the multiplication by 100 allows us to interpret $solvar(\text{rHDF}^{*},g)$ as the percentage of its maximum value.
If solution variability in equation (6) is 0, then there is a single, unique solution to the reconstruction problem in geounit $g$, and the $\text{rHDF}^{*}$ records in $g$ must exactly match the HDF for the variables present in records of both data sets and under the schema used to create $\text{rHDF}^{*}$. When solution variability in equation (6) is 100, there is at least one other reconstruction $\text{rHDF}^{\prime}$ that has no records in common with $\text{rHDF}^{*}$. In this case, $\text{rHDF}^{*}$ may be a poor reconstruction of HDF, and any agreement between the two may be due to happenstance. In general, $solvar$ is an upper bound on the percentage of records in $\text{rHDF}^{*}$ that could differ from any other possible reconstruction, including the actual HDF. For example, if the solution variability is 10, then at most one-tenth of the records (i.e., 10%) in $\text{rHDF}^{*}$ could differ from the actual HDF. Note that “differ” means that at least one attribute of the record is incorrect; it does not necessarily mean all attributes in the record were wrong.
@@
Computing $solvar(\text{rHDF}^{*},b)$ for each block $b$ is manageable using GurobiTM, but computing solvar for larger regions $g$ can be very computationally expensive. Let $\operatorname{Blocks}(g)$ denote the set of blocks composing region $g$ and, as usual, $N_{b}$ and $N_{g}$ refer to the population counts in block $b$ and region $g$, respecitvely. Given $*solvar(\text{rHDF}^{*}$, $b)$ for each $b\in\operatorname{Blocks}(g)$, we define cumulative solution variability, cumsolvar, as:
@@
[…]
4.6. Solvar Implementation Details
The integer program that determines solution variability in geography $g$ is computationally more expensive to solve than solving for an initial $\text{rHDF}^{*}$. For an attacker who wants to target a small number of blocks or tracts, that is not much of an issue. However, the data curator needs to compute the solution variability for all block and tracts. We found that computing a solution to equation (6), in which the feasible $\text{rHDF}^{\prime}$ is required to be consistent with the block and tract level tables, is too expensive to perform for the entire nation. Hence, we relaxed the optimization problem to make it easier to solve while still allowing us to bound the solution variability for an attacker who performs a reconstruction using block and tract-level tables. While concerted research on this problem class is likely to yield faster solutions, we found that the following changes made the problem easier to solve. First, we use the 38-bin age groupings (feature agebin) that appear in the block-level tables.

### v2

@@ 4. Reconstruction Methodology
We define database reconstruction as any attempt to re-create the record-level image of the database from which a set of published query results or tabulations were originally calculated; in this case, that is the confidential HDF. [^17] See garfinkel2018a for a longer discussion of how to understand database reconstruction. Database reconstruction attempts to reverse-engineer the confidential HDF records that were the input data used in a tabulation system with the goal of making the reconstruction as close as possible to these confidential data. We note that the reconstruction described here is not the most powerful feasible reconstruction. We used only a subset of the SF1 tables, we made no attempt to reconstruct households or characteristics of the householder (adult respondent providing the census data on CUF), and we did not use statistical modeling to improve the reconstruction (e.g., in blocks where multiple different reconstructions were consistent with the published tables, we did not use statistical methods to identify the most likely reconstruction). Even without these enhancements, we show that the disclosure avoidance methods of the 2010 Census are highly susceptible to attack.
@@ 4.2. A simplified algebraic representation of the reconstruction problem
where $\mathbf{c}$ is $K\times 1$ column vector of the $K$ statistics extracted from SF1 for a given reconstruction, $\mathbf{A}$ is the $K\times|\chi|$ matrix for computing those SF1 tabulations from a $|\chi|\times 1$ contingency table vector $\mathbf{x}$, and $\mathbb{Z}^{+}$ is the set of non-negative integers. Each row of $\mathbf{A}$ and corresponding component of $\mathbf{c}$, therefore, represent the formula for a single statistic and its realized value in SF1, respectively.
There are fewer statistics published per block than points in the sample space ($|\chi|>K$), meaning that a unique solution to Equation 1 is not always guaranteed. However, many low-population blocks have published tables that contain a large number of entries whose values are zero. This eliminates many candidate solutions, and often there is only one unique non-negative, integer-valued solution to Equation 1. There is always at least one solution because SF1 was tabulated from a single real database with the schema encoded in $\chi$ (i.e., the HDF). The commercial software GurobiTM has a robust toolkit for solving IPs and related mixed integer linear programs (MILP) to produce solutions to problems such as these. See Section 4.3 for details. When multiple solutions exist, GurobiTM will pick one. In principle, statistical modeling could be used to select among candidate solutions to improve reconstruction quality, but we did not do this.
Given a solution $\hat{x}$ to Equation 1, another question of interest is how different $\hat{x}$ could be from other possible solutions. If $\hat{x}$ is the only solution to Equation 1, then it represents an exact reconstruction of the microdata used to tabulate SF1, namely HDF, with certainty. Moreover, if the uniqueness of the solution can be determined from the published inputs to Equation 1, then an attacker knows that those reconstructed records were present in HDF with certainty. To examine how often the solution $\hat{x}$ was strongly constrained, we designed an algorithm for measuring solution variability by building a second IP model. Its goal is to find a second solution, $\tilde{x}$, that maximizes the $\text{L}_{1}$ distance to $\hat{x}$. This allows an attacker, using only public information, to determine the maximum number of records in the reconstruction that could be incorrect. This problem is described in Section 4.5.
@@ 4.3. The integer programming version of the reconstruction model
[…]
@@
AGE =$=\{0,1,2,\dots,110\}$
SEX =$=\{\text{Male}=0,\text{Female}=1\}.$
A potential record corresponds to a setting of the block $b$ along with settings for the demographic attributes (w, bl, aian, asian, nhopi, sor, hisp, a, s). Here the lowercase italic letters index the permitted values of the uppercase feature set of the same name (e.g., $aian$ represents a value of the feature AIAN). We create one variable for the IP for each potential record. Since we do not know, a priori, how many records exist with the same demographic type and block, we must create multiple variables $\mathcal{B}_{(i,b,w,bl,aian,asian,nhopi,sor,hisp,a,s)}$ for $i=0,1,2,\dots$. We use Table P12 from SF1 (age group by sex for each census block) to obtain an upper bound on the index $i$ for each block and combination of demographics. For example, for the demographic type of a 22-year-old Asian Hispanic female in a block $b$, suppose Table P12 for that block indicates that there are 50 females in the age group 22-24 (See Table 4 for the age groupings in P12). This is an upper bound on the number of potential records for 22-year-old Asian Hispanic females in the block. Thus, we create the 50 variables $\mathcal{B}_{(i,b,w=0,bl=0,aian=0,asian=1,nhopi=0,sor=0,hisp=1,a=22,s=1)}$ for $i=0,\dots,49$. These variables are binary, with a value of 1 indicating the presence of the potential record in a candidate microdata reconstruction. Once the binary variables are assigned values, the summation
@@
  • For every block $b$, age $a$, sex $s$, and value of the race and ethnicity variables w, bl, aian, asian, nhopi, sor, hisp, we define the binary variables:
@@
where the summation over $i$ uses the upper bound on the number of $\mathcal{B}_{(\cdot)}$ variables in Equation 3 (i.e., $i=0,\dots,c^{\text{P12}}_{b,\text{AGEBIN$\_$P12}(a),s}-1$).
@@
We use the notational shorthand $\mathcal{T}_{t}[tabname,\bigcup_{b^{\prime}\in\text{B}_{t}}\text{rHDF}_{b,t}(b^{\prime})]=c_{t}^{tabname}$ to represent all such constraints created by SF1 tract-level table $tabname$ (e.g., PCT12I) for tract $t$. Specifically, the notation $\text{rHDF}_{b,t}(b^{\prime})$ refers to the optimization variables for records in a block $b^{\prime}$ when performing the $\text{rHDF}_{b,t}$ reconstruction, the notation $\bigcup_{b^{\prime}\in\text{B}_{t}}\text{rHDF}_{b,t}(b^{\prime})$ indicates that we form the constraints using the optimization variables for all blocks $b^{\prime}$ that belong inside tract $t$ (i.e., $\forall b^{\prime}\in\text{B}_{t}$). For example, in the case of SF1 Table PCT12I, this means an application of Equation 4.3 for each age bin $z$ and sex $s$ for the record variables in the tract. The tract-level tables used are listed in Panel B of Table 2. Similarly, we use the shorthand $\mathcal{T}_{b}[tabname,\text{rHDF}_{b,t}(b^{\prime})]=c_{b^{\prime}}^{tabname}$ for the block-level constraints in block $b^{\prime}$ created by SF1 block-level table $tabname$ (Panel A of Table 2).
@@
The objective function “max 0” indicates that any feasible solution that satisfies the constraints can be returned (i.e., if there are multiple candidate solutions, no statistical modeling or maximum entropy is used, and GurobiTM picks a solution arbitrarily). The optimization variables are the $\mathcal{B}_{(\cdot)}$ defined in Equation 3. Once we have a feasible solution, for every $\mathcal{B}_{(\cdot)}$ that is set to 1, we add a corresponding record for that block and demographic type to the reconstructed dataset. For $\text{rHDF}_{b,t}$, we also recode the feature age into agebin because our matching algorithms require access to both variables (see Section 5). Thus, upon completion, the IP from Equation 4 yields a reconstructed version of the HDF, called $\text{rHDF}_{b,t}$, that contains one record for each person in the 2010 Census with the features indicated in row $\text{rHDF}_{b,t}$ of Table 1.191919 Since SF1 always groups all persons 100 years of age or older into the bins: “100-104 years”, “105-109 years”, and “110 years and over” or uses even coarser age groups, the IP can never resolve the feature age more precisely than these bins. The IP still has variables for individual ages 100, …, 110 and so Gurobi will arbitrarily choose a specific age within those age bins. This means that there is inherent solution variability in the ages of the oldest sub-populations.
Next, we present the IP for the reconstruction $\text{rHDF}_{b}$, which uses only the SF1 block-level tables shown in Panel A of Table 2. Block-level tables use several age binning schemes, but their intersection is not exact age. Instead, their intersection is the age grouping shown in Table 5, which has 38 age bins. Since this is the most fine-grained age resolution that can possibly be obtained from block-level tables, $\text{rHDF}_{b}$ contains age reconstructed up to this 38-age binning (the feature that represents this age grouping is called agebin). To simplify the implementation, the IP uses the $\mathcal{B}_{(\cdot)}$ optimization variables from Equation 3. This means that the solution provides single-year ages, but after the reconstructed $\text{rHDF}_{b}$ microdata are created, we recode age into agebin as defined in Table 5. The IP for the block-level reconstruction is similar to the tract-level reconstruction. For each block $b^{\prime}$, solve
@@
4.4. Reconstruction implementation details
The IP formulation of the reconstruction problem is mathematically equivalent to the simplified model in Section 4.1. The only difference is that the optimization variables $\mathcal{B}_{(\cdot)}$ in the IP implemented in our code correspond to potential records, while the vector $\mathbf{x}$ in Equation 1 contains counts from the fully saturated contingency table defined on the sample space $\chi$. The code in the replication archive implements the IP in Equations 4 and 5, and these equations are useful for reading the software implementation in the replication archive, whereas the histogram representation in Equation 1 is useful for understanding the high-level structure of the problem. For reconstruction outcomes, only run-time, not the space of feasible solutions, is affected by the choice of which representation to implement. It is not obvious a priori which representation should solve more quickly. We have direct experience solving the IP representation, and we found it to consistently solve very quickly at default GurobiTM settings for the set of tables we used in the reconstruction. Equation 1 is succinct, easy to represent in any matrix programming language that implements sparse matrix storage and MILP solvers, and yields a model with considerably fewer variables. On the other hand, the IP representation, while it produces less succinct models, uses only binary variables in the solution set, rather than non-negative integers, and binary variables are usually processed more efficiently in modern MILP solvers. Our discussion switches between the two representations to permit clarity of expression (Equation 1) versus fidelity to the details of our implemented reconstruction code (Equations 4 and 5).
@@ 4.5. Solution variability
where feasible solutions use Equation 4 when $g$ is a tract or Equation 5 when $g$ is a block. Note that the multiplication by 100 allows us to interpret $solvar(\text{rHDF}^{*},g)$ as the percentage of its maximum value.
If solution variability in Equation 6 is zero, then there is a single, unique solution to the reconstruction problem in geounit $g$, and the $\text{rHDF}^{*}$ records in $g$ must exactly match the HDF for the variables present in records of both data sets and under the schema used to create $\text{rHDF}^{*}$. When solution variability in Equation 6 is 100, there is at least one other reconstruction $\text{rHDF}^{\prime}$ that has no records in common with $\text{rHDF}^{*}$. In this case, $\text{rHDF}^{*}$ may be a poor reconstruction of HDF, and any agreement between the two may be due to happenstance. In general, $solvar$ is an upper bound on the percentage of records in $\text{rHDF}^{*}$ that could differ from any other possible reconstruction, including the actual HDF. For example, if the solution variability is 10, then at most one-tenth of the records (i.e., 10%) in $\text{rHDF}^{*}$ could differ from the actual HDF. Note that “differ” means that at least one attribute of the record is incorrect; it does not necessarily mean all attributes in the record were wrong.
@@
Computing $solvar(\text{rHDF}^{*},b)$ for each block $b$ is manageable using GurobiTM, but computing solvar for larger regions $g$ can be very computationally expensive. Let $\operatorname{Blocks}(g)$ denote the set of blocks composing region $g$ and, as usual, $N_{b}$ and $N_{g}$ refer to the population counts in block $b$ and region $g$, respectively. Given $solvar(\text{rHDF}^{*}$, $b)$ for each $b\in\operatorname{Blocks}(g)$, we define cumulative solution variability, cumsolvar, as:
@@
Recalling that $\text{B}_{t}$ is the set of all the blocks in tract $t$, $\text{max}(2cumsolvar(\text{rHDF}^{*}$, $\text{B}_{t}),100)$ is an upper bound on the percentage of records in tract $t$ for which any two solutions to $\text{rHDF}_{b,t}$ can differ. [^21] The analogous claim is not true for $\text{rHDF}_{b}$, because our solution variability models, described in Equations 9 and 10, make use of some tract-level tabulations. By removing tract-level tabulations, $\text{rHDF}_{b}$ is less constrained and so has a larger feasible region. The function cumsolvar is of particular interest for regions $g$ composed of all blocks with $solvar(\text{rHDF}^{*},b)$ no greater than the $q^{\text{th}}$ percentile for different values of $q$ (see Table 6). In particular, we focus our attention on persons who live in the set of blocks with $\textit{cumsolvar}=0$. For these blocks, we know that the attacker is certain that the reconstructed records are in HDF regardless of the method used to compute the reconstruction.
4.6. Solution variability implementation details
The integer program that determines solution variability in geography $g$ is computationally more expensive to solve than solving for an initial $\text{rHDF}^{*}$. For an attacker who wants to target a small number of blocks or tracts, that is not much of an issue. However, the data curator needs to compute the solution variability for all block and tracts. We found that computing a solution to Equation 6, in which the feasible $\text{rHDF}^{\prime}$ is required to be consistent with the block and tract level tables, is too expensive to perform for the entire nation. Hence, we relaxed the optimization problem to make it easier to solve while still allowing us to bound the solution variability for an attacker who performs a reconstruction using block and tract-level tables. While concerted research on this problem class is likely to yield faster solutions, we found that the following changes made the problem easier to solve. First, we use the 38-bin age groupings (feature agebin) that appear in the block-level tables.